OpenStudy (anonymous):

If I have an ordered pair (4,12)<--Like This How would I determine an equation for it?? An equation like y=3x

OpenStudy (lgbasallote):

you need another point

OpenStudy (lgbasallote):

y = 3x is a line...and two points determine a line...so you need another point..make sense?

OpenStudy (anonymous):

Well, I'm stuck on a question that is asking me what equation would be appropriate for the ordered pair (4,12)

OpenStudy (lgbasallote):

what's the whole question?

OpenStudy (anonymous):

Screen shot of my algebra 1 prep cd

OpenStudy (lgbasallote):

well that question is a different thing

OpenStudy (lgbasallote):

to know which one (4,6) is a solution to you have to substitute one by one into the equations (4, 6)

OpenStudy (anonymous):

the y or the x?

OpenStudy (lgbasallote):

for example... the given ordered pair is (4,12) let's check if (4,12) is a solution to y = 1/6 x substitute y = 12 and x = 4 12 = 1/6 (4) 12 = 4/6 12 = 2/3 his is not true...therefore (4,12) is not a solution to y = 1/6 x got it?

OpenStudy (anonymous):

Yes, that's perfect!

OpenStudy (anonymous):

Okay so now I have a different question

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

Some questions I see require me to make an equation for the plots on a graph how do you do that?

OpenStudy (anonymous):

Or they ask that I match the plots to the correct equation(s) Because sometimes there are more then one that can work

OpenStudy (lgbasallote):

you mean create an equation out of a graph?

OpenStudy (lgbasallote):

you know x-intercepts and y-intercepts right?

OpenStudy (anonymous):

I have a little trouble with finding x-intercepts

OpenStudy (lgbasallote):

they are the point where the graph crosses the x-axis

OpenStudy (lgbasallote):



OpenStudy (lgbasallote):

that's the x-intercept